### v1

@@ 4.2 CKSU STPP
The STPP construction CKSU STPP gives rise to a fast algorithm that computes $k=2^{N}$ products of $(m-1)^{N}\times(m-1)^{N}$ matrices, while costing a single convolution in the group $\mathbb{Z}_{m}^{3N}=H^{N}=G$. Since this group is abelian, it has an FFT, which costs $O(m^{3N}\log m^{3N})$. For completeness, we give a full description of the algorithm in Appendix A.
@@
Using other STPP constructions, like the CW tensor [CW87], leads to faster asymptotic algorithms using the same scheme described above.
@@
Suppose the algorithm’s runtime is $C\cdot m^{3N}\log(m^{3N})$ while the naive implementation runtime is $2\cdot 2^{N}\cdot(m-1)^{3N}-2^{N}\cdot(m-1)^{2N}$ (the number of FLOPs). In the table below we give some estimates for the threshold where our algorithm becomes faster, assuming different values of $C$ and $m$. We note that a meticulous calculation of the exact FLOP count for the algorithm would possibly result in bigger constants, which has great effect on the threshold point. Moreover, at these sizes the impracticality of the algorithm stems from the enormous memory usage ($m^{3N}$ for each element in the convolution).
@@
6 Framework For Approximate MM via Convolutions
Suppose we are given input matrices $\left\{\mathbf{A}_{i}\right\}_{i=1}^{k}$ and $\left\{\mathbf{B}_{i}\right\}_{i=1}^{k}$, all of size $n\times n$. Let
$\mathbf{a}=\begin{bmatrix}\mathrm{vec}(\mathbf{A}_{1})\\ \vdots\\ \mathrm{vec}(\mathbf{A}_{k})\end{bmatrix}\quad,\quad\mathbf{b}=\begin{bmatrix}\mathrm{vec}(\mathbf{B}_{1})\\ \vdots\\ \mathrm{vec}(\mathbf{B}_{k})\end{bmatrix}\quad,\quad\mathbf{c}=\begin{bmatrix}\mathrm{vec}(\mathbf{A}_{1}\mathbf{B}_{1})\\ \vdots\\ \mathrm{vec}(\mathbf{A}_{k}\mathbf{B}_{k})\end{bmatrix}\quad\in\mathbb{R}^{kn^{2}}.$
Problem 6.1 (Group Theoretic Approximation).
Given $r\geq 1$ and possibly $\mathbf{a},\mathbf{b}$, find an abelian group $G$ of size $s$ and matrices $\Pi_{A},\Pi_{B},\Pi_{C}\in\mathbb{C}^{s\times kn^{2}}$, which minimize
$\left\lVert\Pi_{C}^{\top}\mathcal{F}^{-1}_{G}\left(\mathcal{F}_{G}(\Pi_{A}\mathbf{a})\odot\mathcal{F}_{G}(\Pi_{B}\mathbf{b})\right)-\mathbf{c}\right\rVert_{2}^{2},$
under the constraint that the tensor $\mathcal{T}_{G,\Pi_{A},\Pi_{B},\Pi_{B}}$ represented by the (Normal Form). has tensor rank $\mathsf{Trk}(\mathcal{T}_{G,\Pi_{A},\Pi_{B},\Pi_{C}})\leq r$, and preferably can be computed in time $\tilde{O}(r)$. Here $\mathcal{F}_{G}$ denotes the Fourier transform for $G$.
Note that the specific presentation of $\mathcal{T}_{G,\Pi_{A},\Pi_{B},\Pi_{C}}$ with the matrices above might have a large inner dimension (i.e., larger than $r$), and still might have another representation (an equivalent normal form) with inner dimension at most $r$. In other words, $s$ is only an upper bound on the tensor rank. As we will see in later sections, this distinction is necessary to unify our approaches under this formulation.
Moreover, it is important to note that the tensor rank does not directly relate to the runtime of the bilinear algorithm. As highlighted before, recursion makes linear transformations done before and after the recursive step essentially “free”, in the sense that the running time is dominated by the tensor rank. This was first presented by [Str69]. In our setting, we wish to avoid recursion for the algorithm to be combinatorial. Under this view, one needs some extra conditions to hold in order for the runtime to match the tensor rank. Examples for this are low rank linear transformations, or structured linear transformations (FFT and the like). Even without the tensor rank directly determining the runtime, the question remains interesting on its own, as low tensor-rank approximations of the matrix multiplication tensor are not well understood (these are “low-accuracy” approximations).
Of course this problem cannot be truly solved using the current tools, so we fall back to heuristics. It is also important to consider two different variations of this problem:
  1. 1.
    Oblivious Setup: $\mathbf{a},\mathbf{b}$ are unknown. A data-dependent solution will use the input compute the matrices $\Pi_{A},\Pi_{B},\Pi_{C}$ online, with running time that does not exceed the bilinear operation running time. A data-agnostic solution will pre-compute the matrices, meaning they are fixed for every $\mathbf{a},\mathbf{b}$ (this will be useful if we have some extra assumptions, like some distribution which $\mathbf{a},\mathbf{b}$ are taken from).
  2. 2.
    Deep Learning Setup: $\mathbf{a}$ is known (we can always switch the roles of $\mathbf{a},\mathbf{b}$ by computing the transposition), and we can pre-compute the matrices $\Pi_{A},\Pi_{B},\Pi_{C}$ based on $\mathbf{a}$ in advance. This is relevant in deep learning, since the model’s weights are known.
Example 6.2.
For the “full” budget $r=n$ and $k=1$, we can take $G=\mathbb{Z}_{n^{3}}$ (using the Vanilla TPP construction) and set $\Pi_{A},\Pi_{B},\Pi_{C}\in\mathbb{C}^{\left|G\right|\times n^{2}}$ to be the indexing matrices. More precisely, for $X\in\left\{A,B,C\right\}$, we note that each row of $\Pi_{X}$ corresponds to an element of the group, while each column of $\Pi_{X}$ corresponds to a position in the matrix, and so by setting $1$ iff this position is embedded into the element matching the row, and 0 otherwise, we obtain $\Pi_{A}\mathbf{a}=a$ and $\Pi_{B}\mathbf{b}=b$ (here $a,b:G\to\mathbb{R}$ are the signals, viewed as vectors). We give these indexing matrices special notation, by writing $I_{X}:=\Pi_{X}$ as defined above (we will reference them later).
@@ 7 Sketch & Solve Approximation
The dominating part is $f(m)=\frac{3\log m-2\log(m-1)}{1+\log(m-1)}$ that approaches $1$ from below as $m\to\infty$ and has a minimum at $m\approx 20$ (for which $f(m)\approx 0.85$). For $m\in[8,480]$, we have $f(m)\leq 0.9$. The second part $g(N)=\frac{\log(3N)}{N(1+\log(m-1))}$ vanishes with $N$. Therefore, we can compute the sketch in time $O(n^{2}r^{\delta+o(1)})$, where $\delta\leq 0.9$ (we can always use small $m$ and take $N$ to be large). The claim the follows from the JL sketch theorem. ∎
This theorem shows that for large enough matrices, we can achieve a better bound on the error, in the same runtime, thus beating the state of the art. In the formulation of 6.1, we can write the Sketch & Solve algorithm as follows:
$\Pi_{A}=I_{A}\vec{S}\quad,\quad\Pi_{B}=I_{B}\vec{S^{\top}}\quad,\quad\Pi_{C}=I_{C},$
where $I_{X}$ are the indexing matrices (see 6.2) and $\vec{S}$ is the natural embedding of the linear sketch $S$, seen as a linear morphism $S:\mathcal{M}_{kn\times kn}(\mathbb{R})\to\mathcal{M}_{kn\times r}(\mathbb{R})$, into the space of linear morphisms $\mathbb{R}^{kn^{2}}\to\mathbb{R}^{knr}$ (and similarly $\vec{S^{\top}}$ is the embedding of $S^{\top}$). Therefore $\Pi_{A},\Pi_{B}$ act as linear sketch “compressions” of the input vector into a lower dimensional group.
Remark 7.3.
The Sketch & Solve scheme is rather unique, in the sense that the dimensions of the $\Pi$ matrices correspond to the underlying tensor rank, which is exactly the group size.
8 Low Degree Embedding
The second approach we will present for solving 6.1 is a non-black box approach. The idea is to embed the matrices into sets which are approximately TPP/STPP sets. For the sake of simplicity of presentation we will focus on TPP, but a similar argument could be made for STPP constructions. Relaxing the TPP condition allows for a smaller group size, and therefore assuming the group has an FFT algorithm, we obtain a fast algorithm to compute an approximate matrix multiplication.
@@ 8.1 Approximate Triple Product Quantity
We then define
@@
Remark 8.3.
The (Approximate Triple Product Quantity). can be related to the notion of Efficacy presented in [AZ23]. In particular, letting $\mathcal{T}_{G,S,T,U}$ denote the tensor obtained by the above approximate embedding scheme and convolution, and using the definition of buckets (C.2) presented in Appendix C, we have
$\mathbf{Eff}(\mathcal{T}_{G,S,T,U})^{2}=\sum_{s\in S,u\in U}\frac{n^{2}}{\left|\mathrm{Bucket}_{G}^{s,u}\left(S,T,U\right)\right|}\quad\text{where}\quad\sum_{s\in S,u\in U}\left|\mathrm{Bucket}_{G}^{s,u}\left(S,T,U\right)\right|=\rho_{n,G}(S,T,U).$
In certain constructions, the ATPQ is balanced, in the sense that buckets have almost constant sizes, which implies the (non-squared) efficacy is $\approx n^{4}/\rho_{n,G}(S,T,U)$. Using the efficacy notion might provide a natural generalization to the ATPQ that easily carries over to the (STPP). case.
@@
Let $G$ be a finite group, $S,A\subset G$, then
$\left|\left\{(x_{1},x_{2},a_{1},a_{2})\in S^{2}\times A^{2}\mid x_{1}x_{2}^{-1}=a_{1}a_{2}^{-1}\right\}\right|\geq\frac{\left|A\right|^{2}\cdot\left|S\right|^{2}}{\left|G\right|}.$
@@
Note that $x_{1}x_{2}^{-1}=a_{1}a_{2}^{-1}$ if and only if $a^{-1}_{1}x_{1}=a_{2}^{-1}x_{2}$, so we can rewrite the LHS by -
$\sum_{g\in G}$ $\left|\left\{(x_{1},x_{2},a_{1},a_{2})\in S^{2}\times A^{2}\mid a_{1}^{-1}x_{1}=g=a_{2}^{-1}x_{2}\right\}\right|$
$=\sum_{g\in G}\left|\left\{(x,a)\in S\times A\mid a^{-1}x=g\right\}\right|^{2}=(*)$
@@
$(*)\geq\frac{1}{\left|G\right|}\left(\sum_{g\in G}\left|\left\{(x,a)\in S\times A\mid a^{-1}x=g\right\}\right|\right)^{2}.$
The proof is complete by noting
$\sum_{g\in G}\left|\left\{(x,a)\in S\times A\mid a^{-1}x=g\right\}\right|=\left|A\right|\cdot\left|S\right|.$
@@
Denote $A=TU^{-1}=T-U=\{t-u\mid t\in T,u\in U\}$, and apply 8.4) on $S,A$ to obtain that the number of solutions to the last equation (equal to $\rho_{n,G}(S,T,U)$) is at least $\left|A\right|^{2}\cdot\left|S\right|^{2}/\left|G\right|$. We finish by noting that $\left|A\right|$ must equal $n^{2}$, or more precisely that $n^{2}=\left|T^{-1}U\right|=\left|TU^{-1}\right|$ by definition of an indexing set, due to the fact that $t_{1}u_{1}^{-1}=t_{2}u_{2}^{-1}$ if and only if $u_{1}t_{1}^{-1}=u_{2}t_{2}^{-1}$ if and only if (using the fact $G$ is abelian) $t_{1}^{-1}u_{1}=t_{2}^{-1}u_{2}$. ∎
For an abelian group with $\left|G\right|=rn^{2}$, we conclude that $\rho_{n}(G)\geq\frac{n^{4}}{r}$. Note that the corollary doesn’t hold for non-abelian groups, because the proof uses commutativity to obtain the expression used in 8.4.
@@ 8.2 PolyForm Algorithm
and a term is summed into $\mathbf{C}$ if and only if it is a solution to the equation $s_{1}s_{2}^{-1}t_{1}t_{2}^{-1}=u_{1}u_{2}^{-1}$. There are $n^{3}$ valid solutions, which are $s_{2}=s_{1},t_{2}=t_{1},u_{2}=u_{1}$, which encode the matrix multiplication, while all other terms are noise terms. In the formulation of 6.1, the matrices $\Pi_{X}$ are just indexing matrices $I_{X}\in\mathcal{M}_{\left|G\right|\times n^{2}}$ which embed into the sets $S^{-1}T,T^{-1}U,S^{-1}U$, and the rank of the resulting tensor is just the tensor of the structure tensor of $G$.
The actual Frobenius error of this algorithm then depends on the squared sum of the coefficients corresponding to these error terms. Since the collisions are pre-determined by the indexing triplet, they might be skewed, meaning that certain positions ($s^{-1}u$) might suffer from relatively large numbers of collisions. Moreover, big error coefficients might end up together under certain orderings of $S,T,U$, but not under different orderings. To mitigate the first effect, we should be mindful of how skewed a certain indexing triplet is. Moreover, to mitigate the latter effect, we can randomly choose the ordering of the indexing sets.
Another randomness we can add to the algorithm is adding structured random signs to $\mathbf{A},\mathbf{B}$. The goal here is to promote error terms cancellations, as we’ll show next. Take $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$ to be independent random vectors $\left\{\pm 1\right\}^{n}$ which are drawn uniformly at random. Define $\tilde{\mathbf{A}}=(\boldsymbol{\alpha}\boldsymbol{\gamma}^{\top})\odot\mathbf{A}$ and $\tilde{\mathbf{B}}=(\boldsymbol{\gamma}\boldsymbol{\beta}^{\top})\odot\mathbf{B}$, where $\odot$ denotes the Hadamard product of matrices. Then $\tilde{\mathbf{A}}_{s,t}=\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\cdot\mathbf{A}_{s,t}$ and $\tilde{\mathbf{B}}_{t,u}=\boldsymbol{\gamma}_{t}\boldsymbol{\beta}_{u}\mathbf{B}_{t,u}$. Since $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$ are independent, we have that
@@
This is exactly what we are interested in – canceling collisions which can be characterized by the condition $t\not=t^{\prime}$.
@@
Note that the algorithm implicitly uses some indexing triplet, thus we define $\mathcal{E}_{G}=\min_{(S,T,U)\in\mathcal{I}(G,n)}\mathcal{E}_{S,T,U}$ as the best algorithm error achieved by the group $G$. The error can be related with the (Approximate Triple Product Quantity)., as described in the next theorem. The proof is given in Appendix C. The general idea of the proof is using Yao’s principle with a distribution over matrices with i.i.d. values in $\left\{\pm 1\right\}$, and showing that in expectation the RPF error is related with $\rho_{n}(G)$.

### v2

@@ 3 Sub-Cubic Matrix Multiplication via Convolutions
The STPP construction CKSU STPP. gives rise to a fast algorithm that computes $k=2^{N}$ products of $(m-1)^{N}\times(m-1)^{N}$ matrices, while costing a single convolution in the group $\mathbb{Z}_{m}^{3N}=H^{N}=G$. Since this group is abelian, it has an FFT, which costs $O(m^{3N}\log m^{3N})$. For completeness, we give a full description of the algorithm in Appendix A.
@@
Using other STPP constructions, like the CW tensor [CW87], may lead to faster asymptotic algorithms using the same scheme described above. However, the basic construction usually shows up in a very high power of the tensor (as in the laser method [AW21]).
@@
Suppose the algorithm’s runtime is $C\cdot m^{3N}\log(m^{3N})$ while the naive implementation runtime is $2\cdot 2^{N}\cdot(m-1)^{3N}-2^{N}\cdot(m-1)^{2N}$ (the number of FLOPs). In the table below we give some estimates for the threshold where our algorithm becomes faster, assuming different values of $C$ and $m$. We note that a meticulous calculation of the exact FLOP count for the algorithm would possibly result in bigger constants, which has great effect on the threshold point. In other words, this is merely a sketch of the runtime asymptotics, which shows how sensitive the threshold point is to the underlying constants. Moreover, at these sizes the impracticality of the algorithm stems from the enormous memory usage ($m^{3N}$ for each element in the convolution).
@@ 4 Sketch & Solve Approximation
The dominating part is $f(m)=\frac{3\log m-2\log(m-1)}{1+\log(m-1)}$ that approaches $1$ from below as $m\to\infty$ and has a minimum at $m\approx 20$ (for which $f(m)\approx 0.85$). For $m\in[8,480]$, we have $f(m)\leq 0.9$. The second part $g(N)=\frac{\log(3N)}{N(1+\log(m-1))}$ vanishes with $N$. Therefore, we can compute the sketch in time $O(n^{2}r^{\delta+o(1)})$, where $\delta\leq 0.9$ (we can always use small $m$ and take $N$ to be large). The claim then follows from the JL sketch theorem. ∎
This theorem shows that for large enough matrices, we can achieve a better bound on the error, in the same runtime, thus beating the state of the art.
5 Approximate Triple Product Polynomials
The second approach we will present is a non-black box approach. The idea is to embed the matrices into sets which are approximately TPP/STPP sets. For the sake of simplicity of presentation we will focus on TPP. Relaxing the TPP condition allows for a smaller group size, and assuming the group has an FFT algorithm, we obtain a fast algorithm to compute an approximate matrix multiplication.
@@ 5.1 Approximate Triple Product Quantity
Equivalently, $\rho_{n,G}(S,T,U)=1_{S}*1_{S^{-1}}*1_{T}*1_{T^{-1}}*1_{U}*1_{U^{-1}}(1_{G})$, where $1_{X}$ is the indicator function of a set $X\subset G$. We also define
@@
Let $G$ be a finite group, $A,B\subset G$, then
$\left|\left\{(a_{1},a_{2},b_{1},b_{2})\in A^{2}\times B^{2}\mid a_{1}a_{2}^{-1}=b_{1}b_{2}^{-1}\right\}\right|\geq\frac{\left|A\right|^{2}\cdot\left|B\right|^{2}}{\left|G\right|}.$
@@
Note that $a_{1}a_{2}^{-1}=b_{1}b_{2}^{-1}$ if and only if $b^{-1}_{1}a_{1}=b_{2}^{-1}a_{2}$, so we can rewrite the LHS by -
$\sum_{g\in G}$ $\left|\left\{(a_{1},a_{2},b_{1},b_{2})\in A^{2}\times B^{2}\mid b_{1}^{-1}a_{1}=g=b_{2}^{-1}a_{2}\right\}\right|=\sum_{g\in G}\left|\left\{(a,b)\in A\times B\mid b^{-1}a=g\right\}\right|^{2}.$ (5.1)
@@
[…]
The proof is complete by noting $\sum_{g\in G}\left|\left\{(a,b)\in A\times B\mid b^{-1}a=g\right\}\right|=\left|A\right|\cdot\left|S\right|$. ∎
@@
Denote $B=TU^{-1}=T-U=\{t-u\mid t\in T,u\in U\}$, and apply 5.3 on $A=S$ and $B$ to obtain that the number of solutions to the last equation (equal to $\rho_{n,G}(S,T,U)$) is at least $\left|S\right|^{2}\cdot\left|T^{-1}U\right|^{2}/\left|G\right|$. We finish by noting that $n^{2}=\left|T^{-1}U\right|=\left|TU^{-1}\right|$ by definition of an indexing set (here we use the fact $G$ is abelian). ∎
For an abelian group with $\left|G\right|=rn^{2}$, we conclude that $\rho_{n}(G)\geq\frac{n^{4}}{r}$. Note that for the inequality to be an equality, we need $\left|\left\{(s,t,u)\in S\times T\times U:s+t-u=g\right\}\right|$ to be a constant for every $g\in G$. For $r<n$ the constant is larger than $1$, and so $(S,T,U)$ cannot be an indexing triplet, at least when $G=\mathbb{Z}_{rn^{2}}$ (for $1_{S}*1_{T}*1_{U^{-1}}$ to be constant, $S,T,U$ must all reside in subgroups of $\mathbb{Z}_{rn^{2}}$, hence they must be arithmetic progressions that can’t avoid each other, meaning that $\left|S^{-1}T\right|\not=n^{2}$, in contradiction with the definition of indexing sets). We conclude that the lower bound is sharp for $r>1$, when $G=\mathbb{Z}_{rn^{2}}$.
Note that the corollary doesn’t hold for non-abelian groups, because the proof uses commutativity to obtain the expression used in 5.3. For non-abelian groups we are able to prove a weaker statement, which depends on the smallest dimension of a non-linear representation of $G$. The proof follows Gowers’ lemma (see [Bre14], Lemma 2.2) and is given in the appendix for completeness.
Lemma 5.5.
Let $d_{\min}=\min\left\{d_{\pi}\mid\pi\in\mathrm{Irr}(G),d_{\pi}>1\right\}$. Then for an indexing triplet $S,T,U\subset G$ we have
$\rho_{n,G}(S,T,U)\geq\frac{(\left|S\right|\left|T\right|\left|U\right|)^{2}}{\left|G\right|}-\left|S\right|\left|T\right|\left|U\right|\cdot\sqrt{\frac{\left|G\right|}{d_{\min}}}.$
Assuming $\left|G\right|\leq rn^{2}$, in order for the lower bound to be non-trivial, we need $d_{\min}>\frac{n^{2}r^{3}}{(n-r)^{2}}\approx r^{3}$. In particular, this lower bound suggests that groups for which $d_{\min}$ is very large, the quantity becomes comparable with that of abelian groups (there is strong mixing).
@@ 5.2 PolyForm Algorithm
and a term is summed into $\mathbf{C}$ if and only if it is a solution to the equation $s_{1}s_{2}^{-1}t_{1}t_{2}^{-1}=u_{1}u_{2}^{-1}$. There are $n^{3}$ valid solutions, which are $s_{2}=s_{1},t_{2}=t_{1},u_{2}=u_{1}$, which encode the matrix multiplication, while all other terms are noise terms.
The actual Frobenius error of this algorithm then depends on the squared sum of the coefficients corresponding to these error terms. Since the collisions are pre-determined by the indexing triplet, they might be skewed, meaning that certain positions ($s^{-1}u$) might suffer from relatively large numbers of collisions. Moreover, big error coefficients might end up together under certain orderings of $S,T,U$, but not under different orderings. Thus we should be mindful of how skewed a certain indexing triplet is. Moreover, to mitigate the latter effect, we can randomly choose the ordering of the indexing sets.
Another randomness we add to the algorithm is structured random signs to $\mathbf{A},\mathbf{B}$. The goal here is to promote error terms cancellations, as we’ll show next. Take $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$ to be independent random vectors $\left\{\pm 1\right\}^{n}$ which are drawn uniformly at random. Define $\tilde{\mathbf{A}}=(\boldsymbol{\alpha}\boldsymbol{\gamma}^{\top})\odot\mathbf{A}$ and $\tilde{\mathbf{B}}=(\boldsymbol{\gamma}\boldsymbol{\beta}^{\top})\odot\mathbf{B}$, where $\odot$ denotes the Hadamard product of matrices. Then $\tilde{\mathbf{A}}_{s,t}=\boldsymbol{\alpha}_{s}\boldsymbol{\gamma}_{t}\cdot\mathbf{A}_{s,t}$ and $\tilde{\mathbf{B}}_{t,u}=\boldsymbol{\gamma}_{t}\boldsymbol{\beta}_{u}\mathbf{B}_{t,u}$. Since $\boldsymbol{\alpha},\boldsymbol{\beta},\boldsymbol{\gamma}$ are independent, we have that
@@
This is exactly what we are interested in – canceling collisions which can be characterized by the condition $t\not=t^{\prime}$. The convolution yields a matrix $\widetilde{\mathbf{C}}$ which we use to obtain the final approximation of $\mathbf{A}\mathbf{B}$, simply by multiplying it with $\boldsymbol{\alpha}^{\top}\boldsymbol{\beta}$ from the left.
@@
Note that the algorithm implicitly uses some indexing triplet, thus we define $\mathcal{E}_{G}=\min_{(S,T,U)\in\mathcal{I}(G,n)}\mathcal{E}_{S,T,U}$ as the best algorithm error achieved by the group $G$.
5.3 Matching lower bound for convolutions in small abelian groups
In this section we prove the lower bound in Theorem 1.4, which generalizes the Cohn and Umans’ [CU03] lower bound of $\left|G\right|\geq n^{3}$ by directly embedding matrix multiplication directly into a convolution.